$SPAD/src/input sersolve.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool sersolve.output
)set message test on
)set message auto off
)clear all
--S 1 of 10
y := operator 'y
--R
--R
--R   (1)  y
--R                                                          Type: BasicOperator
--E 1
--S 2 of 10
eq := D(y x,x) - x*cos(y x) - exp(x)
--R
--R
--R         ,                      x
--R   (2)  y (x) - x cos(y(x)) - %e
--R
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 10
seriesSolve(eq,y,x=0,y(0) = 0)
--R
--I   Compiling function %A with type UnivariateTaylorSeries(Expression(
--R      Integer),x,0) -> UnivariateTaylorSeries(Expression(Integer),x,0)
--R
--R   (3)
--R          2   1  3    1  4    23  5    37  6    61   7    271  8    21617  9
--R     x + x  + - x  - -- x  - --- x  - --- x  + ---- x  + ---- x  + ------ x
--R              6      12      120      360      5040      4032      362880
--R   +
--R      161   10      11
--R     ----- x   + O(x  )
--R     10368
--R                        Type: UnivariateTaylorSeries(Expression(Integer),x,0)
--E 3

)set streams calculate 10
--S 4 of 10
R := EXPR INT
--R
--R
--R   (4)  Expression(Integer)
--R                                                                 Type: Domain
--E 4
--S 5 of 10
uts := UTS(R,'x,0)
--R
--R
--R   (5)  UnivariateTaylorSeries(Expression(Integer),x,0)
--R                                                                 Type: Domain
--E 5
--S 6 of 10
foo: uts -> uts
--R
--R                                                                   Type: Void
--E 6
--S 7 of 10
foo y ==
  xx := monomial(1,1)$uts
  xx * cos(y) + exp(xx)
--R
--R                                                                   Type: Void
--E 7
--S 8 of 10
y := ode1(foo,0)$UTSODE(R,uts)
--R
--R   Compiling function foo with type UnivariateTaylorSeries(Expression(
--R      Integer),x,0) -> UnivariateTaylorSeries(Expression(Integer),x,0)
--R
--R   (8)
--R          2   1  3    1  4    23  5    37  6    61   7    271  8    21617  9
--R     x + x  + - x  - -- x  - --- x  - --- x  + ---- x  + ---- x  + ------ x
--R              6      12      120      360      5040      4032      362880
--R   +
--R      161   10      11
--R     ----- x   + O(x  )
--R     10368
--R                        Type: UnivariateTaylorSeries(Expression(Integer),x,0)
--E 8
--S 9 of 10

x : uts := x
--R
--I   Compiled code for %A has been cleared.
--R
--R   (9)  x
--R                        Type: UnivariateTaylorSeries(Expression(Integer),x,0)
--E 9
--S 10 of 10
x * cos(y) + exp(x)
--R
--R
--R   (10)
--R              1  2   1  3   23  4   37  5    61  6   271  7   21617  8    805  9
--R     1 + 2x + - x  - - x  - -- x  - -- x  + --- x  + --- x  + ----- x  + ---- x
--R              2      3      24      60      720      504      40320      5184
--R   +
--R        841499  10      11
--R     - ------- x   + O(x  )
--R       3628800
--R                        Type: UnivariateTaylorSeries(Expression(Integer),x,0)
--E 10
)spool
)lisp (bye)

———-

References^{[1] nothing}

6